# SyncAnyone: Implicit Disentanglement via Progressive Self-Correction for Lip-Syncing in the wild

### v1

@@ 1 Introduction
Despite their effectiveness, mask-inpainting-based methods [zhang2024personatalk, li2024latentsync, prajwal2020lip] are highly sensitive to mask design. When assigned with small mask regions, the model tends to exploit shortcut solutions by inferring lip movements from contextual cues near the mask boundary, such as chin contours and facial dynamics, thereby bypassing audio conditioning. With large mask regions, critical background and identity-related content are masked out, leading to inaccurate background reconstruction, compromised identity consistency, and failure in challenging cases such as cuts or fast motion.
To mitigate the mask design trade-off, a feasible solution [peng2025omnisyncuniversallipsynchronization] is to adopt mask-free training on collected paired data [10.1007/978-3-030-58589-1_42] with consistent identity, pose, and background but varying lip movements. This help the model to learn both lip editing and visual consistency. However, such data requires highly controlled recording conditions and is difficult to scale, limiting practicality and generalization.
@@ 2.1 GAN-based Methods
Early and mainstream lip synchronization methods predominantly rely on Generative Adversarial Networks (GANs) [goodfellow2020generative, doukas2021headgan, guo2024liveportrait]. Wav2Lip [prajwal2020lip] stands as a landmark work in this field. It pioneered the use of a pre-trained lip-sync discriminator, SyncNet [chung2017out], to supervise the generator’s training, which significantly enhanced the accuracy of the generated lip movements and established a benchmark for subsequent research.
Building on this foundation, subsequent works have introduced improvements from various perspectives. Some methods focus on enhancing generation quality and identity preservation. For instance, DINet [zhang2023dinet] deforms the feature maps of a reference image to generate more natural mouth shapes, thereby better preserving high-frequency details. VideoReTalking [cheng2022videoretalking] decomposes the task into three stages: semantic-guided reenactment, lip synchronization, and identity-aware refinement. StyleSync [guan2023stylesync] employs StyleGAN2 [karras2020analyzing] as its generator backbone, leveraging its powerful generative capabilities to improve visual quality. Most of these methods are arbitrary-subject models, designed to build a general model that does not require retraining for different identities.
Another line of work introduces intermediate representations to guide the generation process. For example, some works first predict facial landmarks [xie2021towards, lu2021live] or 3D-reconstruction-based landmarks [song2022everybody] from audio before performing image-to-image translation. IP-LAP [zhong2023identity] adopts such a two-stage approach. Some methods [guan2024resyncer, zhang2024personatalk] further incorporates 3D mesh priors to guide facial motion, effectively reducing artifacts. Additionally, personalized methods [obama, NVP, adnerf] exist, which can achieve photorealistic results by training on a specific person. However, they suffer from poor generalization and limited applicability.
@@ 2.2 Diffusion-based Methods
More recently, the advent of Diffusion Models has led to a significant leap in performance for audio driven portrait animation [tian2024emo, meng2025mirrorme, li2025infinityhuman, gao2025wan, yang2025infinitetalk], which makes it feasible to perform lip-syncing tasks using diffusion models. with these methods typically generating results of higher resolution and richer detail. They can be broadly categorized into end-to-end and two-stage architectures.
End-to-end models directly generate target video frames conditioned on audio. LatentSync [li2024latentsync] is an end-to-end framework based on an audio-conditioned Latent Diffusion Model (LDM), which requires no intermediate motion representation. SayAnything [ma2025sayanything] and Diff2Lip [mukhopadhyay2024diff2lip] follow a similar idea, performing audio-conditioned denoising generation directly in the latent or pixel space.
Two-stage models decompose the task into audio-to-motion and motion-to-image steps. For example, MyTalk [yu2024make] uses a diffusion model in the first stage to convert audio to motion and a VAE in the second stage for image generation. StyleSync (diffusion version) [zhong2024style] employs a Transformer in its first stage and a diffusion model in its second. DiffDub [liu2024diffdub] utilizes a diffusion autoencoder to convert masked images into semantic latent codes, which are then used alongside audio to condition the final image generation.
To break through the limitations of this paradigm, recent work has begun to explore more universal solutions. For instance, OmniSync [peng2025omnisyncuniversallipsynchronization] pioneered a mask-free framework. It employs a Diffusion Transformer (DiT) to directly edit video frames, aiming to move beyond the dependency on facial priors and extend applicability to arbitrary AI-generated content (AIGC).
@@ 3.1 Preliminaries
Flow Matching. We adopt Flow Matching [lipman2022flow] as our core generative paradigm. It directly learns a deterministic vector field $v_{t}$ that transports a prior distribution (e.g., $\mathcal{N}(0,I)$) to the data distribution. The model, parameterized by $\theta$ and conditioned on a context $c$, is trained by regressing this vector field via a simple L2 objective:
@@ 3.2 Stage 1: Lip Sync via Mask Inpainting
Input Channel Specialization and Training. We specialize the I2V input channels for this task. The conditioning channel $Y$ contains the condition frames $X_{cond}$ with their mouth regions masked out, while the mask channel $M$ is used to differentiate between remain and generate regions. For the remaining region in frames, the mask is set to all-ones, indicating they are fully known. For change region in frames, we apply a large rectangular mask over the lower face, setting these regions to zero to signal for inpainting. The model is then trained end-to-end within the Flow Matching paradigm, conditioned on both the condition frames and the target audio $A_{tgt}$. The objective is to predict the vector field that reconstructs the unmasked portions of $X_{tgt}$ and synthesizes a new mouth in the masked region that corresponds to $A_{tgt}$.
Formally, given a sequence of condition frames $X_{cond}=\{x_{1},\dots,x_{k}\}$, mask $M=\{m_{1},\dots,m_{k}\}$, and a target audio sequence $A_{tgt}=\{a_{k+1},\dots,a_{N}\}$, the model generates the target frames ${X^{\prime}}_{tgt}=\{{x^{\prime}}_{k+1},\dots,{x^{\prime}}_{N}\}$:
@@
Multi-Reference Temporal Modeling. A key aspect of our Stage 1 design is a multi-reference temporal modeling strategy to capture person-specific identity (ID) and dynamic speaking styles. Instead of using a single reference image, we provide the model with a sequence of reference frames. During training, we randomly sample two non-contiguous and non-overlapping segments from a single video clip. One is used as the reference sequence, $X_{ref}=\{x_{1},\dots,x_{k}\}$, and the other as the target sequence to be generate, $X_{tgt}=\{x_{k+1},\dots,x_{N}\}$. By concatenating the reference and target sequences temporally, the self-attention mechanism naturally propagates features from the reference frames to the generation region. This enables the target sequence to draw upon the rich contextual information provided by the reference sequence. During inference, the entire original video clip serves as the reference sequence to generate the new, dubbed clip.
This design offers two significant advantages. First, the abundance of reference frames provides rich and stable identity information, such as lip color, teeth structure, and surrounding skin texture, leading to superior ID consistency. Second, and more importantly, by observing a temporal sequence of the speaker’s natural movements, the model learns their unique speaking modality—such as habitual mouth aperture, speed, and patterns of lip shape transitions—resulting in more personalized and realistic animations.
@@
Audio Feature Injection. To provide a precise driving signal for lip synthesis, the audio condition $A_{tgt}$ is processed and injected into the DiT backbone with careful temporal alignment. Specifically, we first employ a pre-trained Wav2Vec 2.0 model to extract high-level, information-dense feature representations from the raw audio waveform. Inspired by MirrorMe [meng2025mirrorme], we use an causal audio encoder with the same temporal structure of the video vae to align these audio features frame-by-frame with the compressed video latents. The features are then deeply integrated into each layer of the DiT network via cross-attention mechanisms. This allows the model to dynamically attend to the most relevant audio cues at each generation step, ensuring that the synthesized lip movements are precisely synchronized with the nuances of the speech.
@@
Few-Step Generation via DMD. A fully trained, multi-step generative model like $G_{inp}$ is too slow for online data synthesis required in Stage 2. To address this, we distill $G_{inp}$ into a highly efficient, few-step student model, $G^{\prime}_{inp}$, using Distribution Matching Distillation (DMD) [yin2024onestepdiffusiondistributionmatching]. DMD is a state-of-the-art technique that converts a pre-trained flow-based model into a generator that requires only a handful of inference steps (e.g., 1-4) while preserving high generation quality. This distilled model $G^{\prime}_{inp}$ serves as the core synthesis engine for our pseudo-paired data pipeline, making the entire PSC framework computationally feasible.
3.3 Stage 2: Mask-Free Correction
The key innovation of our work lies in Stage 2, where we address the fundamental limitations of explicit spatial guidance. While the Stage 1 model, $G_{inp}$, is robust, its reliance on a pre-defined mask and multi-frame references introduces a risk of background artifacts in motion-heavy scenes (as shown in Fig. 5). More fundamentally, any mask-based approach faces an irresolvable trade-off: a mask that is too small will constrain lip expressiveness and lead to stiff results, while a mask that is too large risks corrupting background details.
To overcome these challenges, we propose a final, efficient, and completely mask-free model, $G_{free}$, that learns to perform implicit disentanglement. The core idea is to train a model that can autonomously deduce which regions of an image to preserve (background and identity) and which to modify (the mouth), conditioned on the target audio and source frames. This is achieved through a novel pseudo-paired data generation pipeline.
Enabling Implicit Disentanglement via Pseudo-Paired Supervision. To teach the model this complex disentanglement task without any spatial masks, we require a unique supervisory signal. We generate this signal by creating a large-scale dataset of $(x_{gt},x_{gen})$ pairs where the background is identical, and only the mouth region differs. The pipeline is detailed in Algorithm 1. We first use a distilled, few-step version of our Stage 1 model ($G^{\prime}_{inp}$) to efficiently generate an initial, mouth-altered frame $x^{\prime}_{gen}$. Then, using a pre-trained face parser, we meticulously correct the synthesized frame by compositing the pristine background from the original frame $x_{gt}$ with the generated face from $x^{\prime}_{gen}$. This process yields a final synthetic frame, $x_{gen}$, which, when paired with the original $x_{gt}$, forms a pseudo-paired training example for our mask-free model.
@@ 3.3 Stage 2: Mask-Free Correction
1:Input: Source video frame $x_{gt}$, random audio $a_{rand}$, distilled model $G^{\prime}_{inp}$, face parser $\mathcal{F}_{parse}$.
2:Output: A pseudo-pair $(x_{gt},x_{gen})$.
3:// Step 1: Generate initial synthetic frame with altered mouth
[…]
5:// Step 2: Compute union mask to robustly cover all facial regions
[…]
9:// Step 3: Composite final synthetic frame with pristine background
[…]
11:return $(x_{gt},x_{gen})$
Learning to Disentangle via a Restoration Objective. With this pseudo-paired dataset, we train our mask-free model, $G_{free}$, using a novel setup designed to force implicit disentanglement through a restoration task. The model is conditioned on the synthetic frame $x_{gen}$ (via the $Y$ channel) and the original audio $A_{gt}$ that corresponds to the ground-truth frame $x_{gt}$. The training objective is to reconstruct the original ground-truth frame $x_{gt}$.
This setup creates a powerful and informative training conflict. The model receives a conditioning frame ($x_{gen}$) that contains a pristine background but an inconsistent (synthesized) mouth. Simultaneously, it receives the audio signal ($A_{gt}$) that perfectly matches the correct mouth in the target ground-truth frame ($x_{gt}$). To successfully reconstruct $x_{gt}$, the model is compelled to learn a highly sophisticated mapping:
  1. 1.
    Preserve the reliable information: It must learn to trust and copy the background and identity cues from the conditioning frame $x_{gen}$.
  2. 2.
    Discard and replace the unreliable information: It must learn to identify and ignore the synthesized mouth region in $x_{gen}$, recognizing it as inconsistent with the provided audio $A_{gt}$.
  3. 3.
    Synthesize from audio: It must then synthesize the correct, natural mouth shape based solely on the audio signal $A_{gt}$.

### v3

@@ 1 Introduction
Despite their effectiveness, mask-inpainting-based methods (Zhang et al., 2024a; Li et al., 2024; Prajwal et al., 2020) are highly sensitive to mask design. When assigned with small mask regions, the model tends to exploit shortcut solutions by inferring lip movements from contextual cues near the mask boundary, such as chin contours and facial dynamics, thereby bypassing audio conditioning. With large mask regions, critical background and identity-related content are masked out, leading to inaccurate background reconstruction, compromised identity consistency, and failure in challenging cases such as cuts or fast motion.
To mitigate the mask design trade-off, a feasible solution (Peng et al., 2025) is to adopt mask-free training on collected paired data (Wang et al., 2020) with consistent identity, pose, and background but varying lip movements. This help the model to learn both lip editing and visual consistency. However, such data requires highly controlled recording conditions and is difficult to scale, limiting practicality and generalization.
@@ 2.1 GAN-based Methods
Early and mainstream lip synchronization methods predominantly rely on Generative Adversarial Networks (GANs) (Goodfellow et al., 2020; Doukas et al., 2021; Guo et al., 2024). Wav2Lip (Prajwal et al., 2020) stands as a landmark work in this field. It pioneered the use of a pre-trained lip-sync discriminator, SyncNet (Chung and Zisserman, 2017), to supervise the generator’s training, which significantly enhanced the accuracy of the generated lip movements and established a benchmark for subsequent research.
Building on this foundation, subsequent works have introduced improvements from various perspectives. Some methods focus on enhancing generation quality and identity preservation. For instance, DINet (Zhang et al., 2023) deforms the feature maps of a reference image to generate more natural mouth shapes, thereby better preserving high-frequency details. VideoReTalking (Cheng et al., 2022) decomposes the task into three stages: semantic-guided reenactment, lip synchronization, and identity-aware refinement. StyleSync (Guan et al., 2023) employs StyleGAN2 (Karras et al., 2020) as its generator backbone, leveraging its powerful generative capabilities to improve visual quality. Most of these methods are arbitrary-subject models, designed to build a general model that does not require retraining for different identities.
Another line of work introduces intermediate representations to guide the generation process. For example, some works first predict facial landmarks (Xie et al., 2021; Lu et al., 2021) or 3D-reconstruction-based landmarks (Song et al., 2022) from audio before performing image-to-image translation. IP-LAP (Zhong et al., 2023) adopts such a two-stage approach. Some methods (Guan et al., 2024; Zhang et al., 2024a) further incorporates 3D mesh priors to guide facial motion, effectively reducing artifacts. Additionally, personalized methods (Suwajanakorn et al., 2017; Thies et al., 2020; Guo et al., 2021) exist, which can achieve photorealistic results by training on a specific person. However, they suffer from poor generalization and limited applicability.
@@ 2.2 Diffusion-based Methods
More recently, the advent of Diffusion Models has led to a significant leap in performance for audio driven portrait animation (Tian et al., 2024; Meng et al., 2025; Li et al., 2025; Gao et al., 2025; Yang et al., 2025), which makes it feasible to perform lip-syncing tasks using diffusion models. with these methods typically generating results of higher resolution and richer detail. They can be broadly categorized into end-to-end and two-stage architectures.
End-to-end models directly generate target video frames conditioned on audio. LatentSync (Li et al., 2024) is an end-to-end framework based on an audio-conditioned Latent Diffusion Model (LDM), which requires no intermediate motion representation. SayAnything (Ma et al., 2025) and Diff2Lip (Mukhopadhyay et al., 2024) follow a similar idea, performing audio-conditioned denoising generation directly in the latent or pixel space.
Two-stage models decompose the task into audio-to-motion and motion-to-image steps. For example, MyTalk (Yu et al., 2024) uses a diffusion model in the first stage to convert audio to motion and a VAE in the second stage for image generation. StyleSync (diffusion version) (Zhong et al., 2024) employs a Transformer in its first stage and a diffusion model in its second. DiffDub (Liu et al., 2024) utilizes a diffusion autoencoder to convert masked images into semantic latent codes, which are then used alongside audio to condition the final image generation.
To break through the limitations of this paradigm, recent work has begun to explore more universal solutions. For instance, OmniSync (Peng et al., 2025) pioneered a mask-free framework. It employs a Diffusion Transformer (DiT) to directly edit video frames, aiming to move beyond the dependency on facial priors and extend applicability to arbitrary AI-generated content (AIGC).
@@ 3.1 Preliminaries
Flow Matching. We adopt Flow Matching (Lipman et al., 2022) as our core generative paradigm. It directly learns a deterministic vector field $v_{t}$ that transports a prior distribution (e.g., $\mathcal{N}(0,I)$) to the data distribution. The model, parameterized by $\theta$ and conditioned on a context $c$, is trained by regressing this vector field via a simple L2 objective:
@@ 3.2 Stage 1: Lip Sync via Mask Inpainting
Input Channel Specialization and Training.
We adapt the I2V-inspired input structure for the mask-inpainting task by specializing its parallel channels. Let $X_{tgt}=\{x_{1},\dots,x_{k}\}$ denote a sequence of ground-truth target frames sampled from the dataset, which corresponds to the target audio $A_{tgt}$. To guide the synthesis, we define a spatial mask $M\in\{0,1\}^{H\times W}$, where a value of 1 indicates the background regions to be preserved, and 0 indicates the mouth region to be inpainted (typically a large rectangular area over the lower face). The conditioning frames $X_{cond}$ are formally derived by masking the ground-truth frames:
$X_{cond}=X_{tgt}\odot M,$ (3)
where $\odot$ denotes the element-wise (Hadamard) product.
The model is trained end-to-end within the Flow Matching paradigm, taking $X_{cond}$, $M$, and $A_{tgt}$ as joint conditions. The training objective is to predict a vector field that reconstructs the original $X_{tgt}$ by synthesizing high-fidelity lip movements within the masked regions that are synchronized with $A_{tgt}$. Formally, the generation of the predicted frames $X^{\prime}_{tgt}$ is expressed as:
@@
During training, the network minimizes the reconstruction error between the generated $X^{\prime}_{tgt}$ and the ground-truth $X_{tgt}$, forcing the model to learn the spatial context from the unmasked areas while following the phonetic cues from the audio.
Multi-Reference Temporal Modeling. To capture person-specific identity and dynamic speaking styles, Stage 1 employs a multi-reference temporal modeling strategy. Unlike methods relying on a single reference image, we provide the model with a sequence of frames to exploit long-range spatiotemporal dependencies.
During training, we process a total of $N$ frames, where the first $k$ frames serve as the reference sequence $X_{ref}=\{x_{1},\dots,x_{k}\}$ and the remaining NkN-k frames constitute the target sequence $X_{tgt}=\{x_{k+1},\dots,x_{N}\}$. These two sequences are sampled from non-contiguous and non-overlapping segments of the same video. The conditioning sequence $X_{cond}$ is then formally constructed by temporally concatenating the reference frames with the masked target frames:
$X_{cond}=[X_{ref}\ ;\ X_{tgt}\odot M],$ (5)
where $M$ masks out the mouth region only for the target frames $\{x_{k+1},\dots,x_{N}\}$, while the reference frames $X_{ref}$ remain fully intact to provide dense identity priors. By feeding this joint sequence into the DiT, the global self-attention mechanism naturally propagates high-fidelity texture and habitual motion features from $X_{ref}$ to the generation regions in $X_{tgt}$. To ensure the model focuses on lip synthesis, the training loss is calculated exclusively on the target sequence:
$\mathcal{L}=\mathbb{E}\|X^{\prime}_{k+1:N}-X_{k+1:N}\|^{2}_{2}.$ (6)
During inference, given a source video $X_{src}$ to be dubbed, we use $X_{src}$ in a dual role: it acts as $X_{ref}$ to provide identity context and simultaneously serves as the masked target sequence $X_{tgt}\odot M$ to provide the driving pose and background. This ensures the generated clip preserves the original head motion and environmental consistency. For visual clarity in Fig. 2 (Stage 1), we simplify the depiction of noisy latents and mask channels; however, the network input strictly follows the channel-wise concatenation described in Eq. 2.
@@
Audio Feature Injection. To provide a precise driving signal for lip synthesis, the audio condition $A_{tgt}$ is processed and injected into the DiT backbone with careful temporal alignment. Specifically, we first employ a pre-trained Wav2Vec 2.0 model to extract high-level, information-dense feature representations from the raw audio waveform. Inspired by MirrorMe (Meng et al., 2025), we use an causal audio encoder with the same temporal structure of the video vae to align these audio features frame-by-frame with the compressed video latents. The features are then deeply integrated into each layer of the DiT network via cross-attention mechanisms. This allows the model to dynamically attend to the most relevant audio cues at each generation step, ensuring that the synthesized lip movements are precisely synchronized with the nuances of the speech.
@@
Few-Step Generation via DMD. A fully trained, multi-step generative model like $G_{inp}$ is too slow for online data synthesis required in Stage 2. To address this, we distill $G_{inp}$ into a highly efficient, few-step student model, $G^{\prime}_{inp}$, using Distribution Matching Distillation (DMD) (Yin et al., 2024). DMD is a state-of-the-art technique that converts a pre-trained flow-based model into a generator that requires only a handful of inference steps (e.g., 1-4) while preserving high generation quality. This distilled model $G^{\prime}_{inp}$ serves as the core synthesis engine for our pseudo-paired data pipeline, making the entire PSC framework computationally feasible.
3.3 Stage 2: Mask-Free Correction and Refinement
While the Stage 1 model ($G_{inp}$) provides robust motion priors, its reliance on explicit spatial masks and multi-frame concatenation introduces an inherent mask design dilemma: a restrictive mask limits lip expressiveness, whereas an expansive mask risks corrupting identity and background details (as shown in Fig. 5). To resolve this, Stage 2 introduces a mask-free tuning paradigm that transitions the model from explicit spatial guidance to Implicit Disentanglement. This stage optimizes a highly efficient generator, $G_{free}$, that learns to autonomously identify and modify only the mouth-related pixels by resolving cross-modal inconsistencies.
Implicit Disentanglement via Restoration-based Supervision. To train $G_{free}$ without spatial masks, we require a supervisory signal that defines what to preserve and what to edit. We generate this signal via a novel pseudo-paired data pipeline (Algorithm 1). For a ground-truth frame $x_{gt}$ and its original audio $A_{gt}$, we first synthesize an initial mouth-altered frame $x^{\prime}_{gen}$ using the distilled Stage 1 model ($G^{\prime}_{inp}$) conditioned on a random audio $A_{rand}$. To ensure the background remains pixel-perfect, we employ a face parser (Yakhyokhuja, 2024) to compute a union mask $M_{face}$, compositing the generated face onto the pristine background of $x_{gt}$. This yields a pseudo-pair where $x_{gen}$ (condition) and $x_{gt}$ (target) share an identical background but possess conflicting lip shapes.
@@ 3.3 Stage 2: Mask-Free Correction and Refinement
Input: Source video frame $x_{gt}$, random audio $a_{rand}$, distilled model $G^{\prime}_{inp}$, face parser $\mathcal{F}_{parse}$.
Output: A pseudo-pair $(x_{gt},x_{gen})$.
// Step 1: Generate initial synthetic frame with altered mouth
$x^{\prime}_{gen}\leftarrow G^{\prime}_{inp}(x_{gt},a_{rand})$
// Step 2: Compute union mask to robustly cover all facial regions
$M^{\prime}_{gen}\leftarrow\mathcal{F}_{parse}(x^{\prime}_{gen})$ {Mask from generated frame}
$M_{gt}\leftarrow\mathcal{F}_{parse}(x_{gt})$ {Mask from ground-truth frame}
$M_{face}\leftarrow M^{\prime}_{gen}\lor M_{gt}$ {Compute the logical OR (union) of masks}
// Step 3: Composite final synthetic frame with pristine background
$x_{gen}\leftarrow M_{face}\odot x^{\prime}_{gen}+(1-M_{face})\odot x_{gt}$
return $(x_{gt},x_{gen})$
Table 1: Quantitative comparison on HDTF and VFHQ.
Method HTDF VFHQ
FID $\downarrow$ FVD $\downarrow$ CSIM $\uparrow$ LSE-C $\uparrow$ LMD $\downarrow$ FID $\downarrow$ FVD $\downarrow$ CSIM $\uparrow$ LSE-C $\uparrow$ LMD $\downarrow$
Wav2Lip (Prajwal et al., 2020) 15.34 562.1 0.841 7.58 9.82 19.12 642.5 0.825 7.42 11.85
VideoRetalking (Cheng et al., 2022) 12.21 392.4 0.772 7.05 8.64 15.38 435.1 0.751 6.94 9.47
IP-LAP (Zhong et al., 2023) 9.82 338.2 0.795 7.14 7.85 11.24 372.6 0.784 7.03 8.66
MuseTalk (Zhang et al., 2024b) 9.12 245.8 0.854 6.75 8.92 10.05 274.3 0.843 6.65 9.82
LatentSync (Li et al., 2024) 8.78 228.3 0.842 8.12 16.85 9.56 256.4 0.835 8.04 17.94
Ours ($1^{st}$ stage) 8.65 220.5 0.858 7.85 7.62 9.35 248.2 0.848 7.72 8.45
Ours ($2^{nd}$ stage) 8.42 205.4 0.865 7.95 7.70 9.12 232.5 0.856 7.84 8.52
Mechanism of the Training Conflict. The core of our implicit disentanglement lies in the restoration objective: we force the model to reconstruct the pristine $x_{gt}$ from the ”noisy” condition $x_{gen}$ under the guidance of the target audio $A_{gt}$. This setup creates a powerful training conflict that compels the network to learn a sophisticated tripartite mapping:
Preserve (Identity/Background): The model must recognize that the background and non-mouth facial regions in $x_{gen}$ are consistent with $x_{gt}$ and should be preserved via a ”copy-and-paste” behavior.
Discard (Inconsistent Motion): It must identify that the mouth region in $x_{gen}$ (driven by $A_{rand}$) contradicts the target audio $A_{gt}$ and should be treated as undesirable noise.
Synthesize (Audio-driven Lip-sync): It must synthesize a new, synchronized mouth shape solely by attending to $A_{gt}$, seamlessly integrating it into the preserved facial context.
Through this conflict-resolution process, the network’s receptive field for modification becomes purely data-driven. The model autonomously learns to disentangle static identity from dynamic motion without any predefined spatial boundaries, effectively solving the mask dilemma.
Robustness to Synthetic Artifacts. A potential concern with pseudo-paired data is the propagation of blending artifacts from the background compositing module. However, our restoration paradigm is inherently self-correcting: because the target $x_{gt}$ is always a pristine, artifact-free original frame, the model learns to map ”synthetic-and-potentially-flawed” inputs to ”real-and-clean” outputs. Consequently, at inference time, when provided with high-quality original video frames as $x_{gen}$, the model remains robust and produces results with lossless background fidelity.
4 Experiment